OpenStudy (i_love_my_nieces):

I NEED HELP WITH 5 QUESTION, I WILL GIVE YOU A MEDAL AND BECOME YOUR FAN! :) 0.o o.0

OpenStudy (i_love_my_nieces):

1.) 7,911,13,... (1 point) Arithmetic Geometric Both Neither

OpenStudy (tkhunny):

Subtract each pair of successive terms. What do you observe? 9 - 7 = ?? 11 - 9 = ?? 13 - 11 = ??

OpenStudy (i_love_my_nieces):

2

OpenStudy (tkhunny):

Every time?

OpenStudy (i_love_my_nieces):

Yeah.

OpenStudy (tkhunny):

Well, then if we ASSUME that this pattern continues indefinitely, this is the definition of an Arithmetic Sequence. Are you willing to assume that?

OpenStudy (i_love_my_nieces):

I don't know so A?

OpenStudy (tkhunny):

You have to be willing to assume the continuation of the pattern. If so, yes. It is A Do you know how to detect a Geometric Sequence?

OpenStudy (i_love_my_nieces):

IS it A I think it is?!?!?!?!!?!?!

OpenStudy (tkhunny):

Didn't we already say that? The same SORT of thing will find Geometric Sequences. Just divide, rather than subtract. If it fails both tests, it is "Other" or "Neither". It CANNOT pass both tests. "Both" is just silly.